# On the comparison of ultradifferentiable weight function settings

### v1

@@ 2.8. Different notions of weight functions
    If $(\omega_{\operatorname{snq}})$ holds then $\omega$ is called strong non-quasianalytic or even strong for short. This terminology is inspired by [5, Def. 1.8].
@@
Next we move to the Beurling-Björk setting.
Definition 2.10.
More precisely, in [3]
@@
    according to [3, Def. 1.2.5] the set $\mathfrak{M}_{0}$ denotes functions satisfying (i)(iv)(i)-(iv).
For the sake of completeness let us mention that also in [23] and [7, Sect. 8] the notion of BB-weights has been considered; in [7] for a comparison of different theories.
More precisely, in [7, 8.4 $(2)$] for weights in the sense of Beurling-Björck all requirements from Definition 2.10 except (formally) continuity have been considered.
Finally, in [23] H.-J. Petzsche and D. Vogt have assumed that
Occasionally, the authors have also considered
Note: $(\delta)$ in [23, Sect. 5] does not mean $(\omega_{4})$ (i.e. the convexity of $\varphi_{\omega}$) and $(\gamma_{1})$ must not be mixed and confused with Petzsche’s condition for weight sequences introduced in [22].
@@ 2.9. Matrix admissible weight functions
Summarizing, assume that $\omega$ is non-decreasing and such that $(\omega_{3})$ holds. Then $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$, so $\omega$ is a weight function in the notion of [12] (and [27], [26]), and by (a)(e)(a)-(e) it follows that each $\mathbf{W}^{(\ell)}$ is a well-defined sequence of non-negative real numbers, each $\mathbf{W}^{(\ell)}$ is log-convex (i.e. $(M.1)$ in [15] is valid), $\lim_{p\rightarrow+\infty}(W^{(\ell)}_{p})^{1/p}=+\infty$ for all $\ell>0$, and finally $W^{(\ell_{1})}_{p}\leq W^{(\ell_{2})}_{p}$ for all $p\in\mathbb{N}$, $0<\ell_{1}\leq\ell_{2}$. Thus basically “non-decreasing” and $(\omega_{3})$ for $\omega$ are sufficient to introduce $\mathcal{M}_{\omega}$ and this explains the notion “matrix admissible”.
If $\omega$ is also normalized, then each $\mathbf{W}^{(\ell)}$ is normalized; i.e. $1=W^{(\ell)}_{0}\leq W^{(\ell)}_{1}$. Note that, when $\omega$ is not normalized, then one can always switch to an equivalent weight $\sigma$ which is normalized and this notion of equivalence transfers in a precise manner to the corresponding associated weight matrices and which ensures the equality of the corresponding (matrix) weighted spaces; we refer to [28, Sect. 6 $(III)$, $(6.6)$].
@@ 3.1. Basic definitions
(3.2) $\mathcal{L}^{\infty}_{\{\mathcal{W}\}}:=\bigcup_{\ell>0}\mathcal{L}^{\infty}_{\omega^{\ell}},\hskip 15.0pt\mathcal{L}^{\infty}_{(\mathcal{W})}:=\bigcap_{\ell>0}\mathcal{L}^{\infty}_{\omega^{\ell}},$
endowed with their natural topologies. Moreover, for any $1\leq p<\infty$ we assume that $\mathcal{W}$ is continuous and for all (parameters) $\ell>0$ let us set
@@
(3.3) $\mathcal{L}^{p}_{\{\mathcal{W}\}}:=\bigcup_{\ell>0}\mathcal{L}^{p}_{\omega^{\ell}},\hskip 15.0pt\mathcal{L}^{p}_{(\mathcal{W})}:=\bigcap_{\ell>0}\mathcal{L}^{p}_{\omega^{\ell}}.$
@@ 3.2. On the unboundedness of weight function matrices
    (3.4) holds if and only if $L^{\infty}=\mathcal{L}^{\infty}_{\{\mathcal{W}\}}$ as sets and/or l.c.v.s. and (3.5) if and only if $L^{\infty}=\mathcal{L}^{\infty}_{(\mathcal{W})}$ as sets and/or l.c.v.s.
@@
    If $\mathcal{W}$ is continuous, then (3.4) holds if and only if $L^{p}=\mathcal{L}^{p}_{\{\mathcal{W}\}}$ and (3.5) if and only if $L^{p}=\mathcal{L}^{p}_{(\mathcal{W})}$ and the equality holds for some/any $1\leq p<\infty$ as sets and/or l.c.v.s.
@@ 3.3. On the non-triviality of weighted $L^{p}$-type spaces
Proof. $(i)$ If (3.5) is valid, then $\mathcal{L}^{\infty}_{(\mathcal{W})}=L^{\infty}\neq\{0\}$ by Theorem 3.3.
@@
(3.9) $\psi^{\infty}_{\mathcal{W}}(x):=e^{-\sigma^{\infty}(x)^{2}}.$
So for any $\ell>0$ and all $x\in\mathbb{R}^{d}$ such that $|x|\in[n,n+1)$ one has $|\psi^{\infty}_{\mathcal{W}}(x)|e^{\omega^{\ell}(x)}=e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{2}}$. Let $\ell\geq\ell_{0}$ be fixed and hence for all $|x|\in[n,n+1)$ with $n\geq\ell$ we have $|\psi^{\infty}_{\mathcal{W}}(x)|e^{\omega^{\ell}(x)}\leq e^{\omega^{\ell}(x)(1-\omega^{\ell}(x))}$ and this last expression tends to 0 as $|x|\rightarrow+\infty$. Thus,
@@
$(ii)$ If (3.5) holds, then $\mathcal{L}^{p}_{(\mathcal{W})}=L^{p}\neq\{0\}$ by Theorem 3.3.
@@
When enlarging $a_{n}$, if necessary, it suffices to take
$b_{n}=\frac{a_{n}\log(\omega^{n}(n+1))+\log(2)}{\log(\omega^{n}(n))}(\geq\max\{1,\frac{\log(\omega^{n}(n+1))+\log(2)}{\log(\omega^{n}(n))}\}).$
@@
$=\sum_{n=0}^{\ell-1}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{a_{n}b_{n}}}dx+\sum_{n=\ell}^{+\infty}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{a_{n}b_{n}}}dx$
$\leq\sum_{n=0}^{\ell-1}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{a_{n}b_{n}}}dx+\sum_{n=\ell}^{+\infty}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{n}(x)-(\omega^{n}(x))^{a_{n}b_{n}}}dx$
$\leq\sum_{n=0}^{\ell-1}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{a_{n}b_{n}}}dx+\sum_{n=\ell}^{+\infty}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}\frac{1}{(1+|x|^{2})^{d+1}}dx<+\infty.$
@@ 4. Characterization of inclusion relations for weighted $L^{p}$-type spaces
Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ and $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be weight function matrices and assume that both matrices satisfy (3.13). Then the following are equivalent:
@@
Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ and $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be weight function matrices and assume that both matrices satisfy (3.14). Then the following are equivalent:
@@ 4.1. Proof of the main characterizing results
Proof. $(i)$ First, by (3.13) we get:
@@
Now, in view of Remark 4.4 we can assume that there exists $0\neq f\in\mathcal{L}^{p}_{(\mathcal{S})}$ and $(i)$ in Proposition 3.6 yields $f_{x_{0}}\in\mathcal{L}^{p}_{(\mathcal{S})}$ for all $x_{0}\in\mathbb{R}^{d}$. The idea is to apply (4.3) to the family $\{f_{x_{0}}:x_{0}\in\mathbb{R}^{d}\}$.
Let $\ell>0$ be given, arbitrary but from now on fixed. Let $\ell^{\prime}>0$ be the index appearing in the estimate in the first step and being related to $\ell$ via (3.13). Then apply (4.3) to $\ell^{\prime}$ and let $\ell^{\prime\prime}$ be the index subject to $\ell^{\prime}$ in this property.
@@
The next result deals with the analogous result for the Roumieu case.
@@ 4.2. Alternative proofs for the Roumieu-type
(4.10) $\exists\;\ell^{\prime}_{0}>\ell_{0}>0:\;\;\;\lim_{|x|\rightarrow+\infty}\omega^{\ell^{\prime}_{0}}(x)-\omega^{\ell_{0}}(x)=+\infty,$
@@
    If (4.9), (4.10) holds for some indices $\ell^{\prime}_{0}>\ell_{0}$, then also for all $\ell^{\prime},\ell$ satisfying $\ell^{\prime}\geq\ell^{\prime}_{0}>\ell_{0}\geq\ell>0$.
@@
    One shall note that (4.9), (4.10) and (4.11) are crucial for the techniques in this section but they can never be satisfied for the simple weight function matrix $\mathcal{W}:=\{\omega\}$. On the other hand, the estimates in (3.13) and (3.14) hold for any sub-additive $\omega$ and hence the main results Theorems 4.1, 4.2 and Proposition 3.6 can be applied to the simple case, too.
@@
$\|\theta^{p}_{\omega^{\ell}}\|^{p}_{p,\omega^{\ell^{\prime}}}=\int_{\mathbb{R}^{d}}e^{\omega^{\ell^{\prime}}(x)-\omega^{\ell}(x)}dx\geq\int_{\mathbb{R}^{d}}e^{\omega^{\ell^{\prime}_{0}}(x)-\omega^{\ell_{0}}(x)}dx=+\infty.$
(However, in general it is not clear if already (4.9) implies this estimate.) ∎
@@
Note: This statement is more general than $(ii)$ in Proposition 4.6 since no stability under translation is required.
@@
Proof. $(iii)\Rightarrow(i)$ follows immediately by definition of the spaces and $(i)\Rightarrow(ii)$ is clear. For $(ii)\Rightarrow(iii)$ follow the proof of Proposition (4.9) and replace $\exists\;\ell^{\prime}>0$” by $\forall\;\ell^{\prime}>0$”. ∎
@@
    (3.14) implies (4.14) and (3.13) implies (4.15) when we set $s:=1$.
@@
thus $\mathcal{S}\{\preceq\}\mathcal{T}$ is verified (between the indices $\ell^{\prime\prime}$ and $\ell_{1}$).
@@
And for $x\in\mathbb{R}^{d}$ with $0\leq|x|<1$ one has $\tau^{\ell^{\prime\prime}_{1}}(x)\leq\tau^{\ell^{\prime\prime}_{1}}(1)\leq\sigma^{\ell}(x)+A$ with $A:=\tau^{\ell^{\prime\prime}_{1}}(1)$. Summarizing, again $\mathcal{S}\{\preceq\}\mathcal{T}$ is verified (between the indices $\ell$ and $\ell^{\prime\prime}_{1}$). ∎
@@
Without the additional assumptions on the weight matrices it is not clear that (4.16) implies $\lim_{|x|\rightarrow+\infty}e^{\tau^{\ell_{1}}(x)-\sigma^{\ell}(x)}=0$ or, what is enough to conclude, $\sup_{x\in\mathbb{R}^{d}}e^{\tau^{\ell_{1}}(x)-\sigma^{\ell}(x)}<+\infty$. This is due to the fact that in general $t\mapsto e^{\tau^{\ell_{1}}(t)-\sigma^{\ell}(t)}$ might behave very irregular: This function can have (infinitely many) “exploding peaks” but being sufficiently small “nearly everywhere on $\mathbb{R}$” and hence (4.16) still can be valid. In view of Remark 4.7 this comment becomes relevant in the simple case.

### v2

@@ 2.8. Comments on different notions of weight functions
    If $(\omega_{\operatorname{snq}})$ holds then $\omega$ is called strong non-quasianalytic or even strong for short. This notion is inspired by [5, Def. 1.8].
@@
Concerning BB-weights in [3]
@@
    according to [3, Def. 1.2.5] the set $\mathfrak{M}_{0}$ denotes functions satisfying (i)(iv)(i)-(iv) in Definition 1.2.
For the sake of completeness let us mention that also in [7, Sect. 8] the notion of BB-weights has been considered for a comparison of different theories. More precisely, in [7, 8.4 $(2)$] for weights in the sense of Beurling-Björck all requirements from Definition 1.2 except (formally) continuity have been assumed.
Finally, concerning PV-weights in [23] (recall Definition 1.3) occasionally the following conditions appear:
  • $(5)$
    $(\omega_{3})$ (abbreviated by $(\gamma_{1})$),
  • $(6)$
    $(\omega_{2})$ (abbreviated by $(\beta_{0})$), and
  • $(7)$
    $(\delta)$ which is
    $\exists\;H\geq 1:\;\;\;\liminf_{t\rightarrow+\infty}\frac{\omega(Ht)}{\omega(t)}>1.$
Thus, $(\delta)$ in [23, Sect. 5] does not mean $(\omega_{4})$ (i.e. the convexity of $\varphi_{\omega}$) and $(\gamma_{1})$ must not be mixed and confused with Petzsche’s condition for weight sequences introduced in [22]. Since $\omega$ is non-decreasing in $(\delta)$ we necessarily have to choose $H>1$ and, indeed, via (2.11) and (2.12) condition $(\delta)$ does precisely mean $\overline{\gamma}(\omega)<+\infty$; i.e. $(\omega_{6})$ for $\omega$. And this observation (for the associated weight function case $\omega=\omega_{\mathbf{M}}$) is consistent with [23, Lemma 5.3] and the known characterization of $(\omega_{6})$ in terms of $\mathbf{M}$; see e.g. [30, Thm. 3.1] and the references there.
@@ 2.9. Matrix admissible weight functions
In [28] and [24] we have developed the idea that to each BMT-weight function $\omega$ we can associate the weight matrix
(2.13) $\mathcal{M}_{\omega}:=\{\mathbf{W}^{(\ell)}:\ell>0\},\hskip 15.0ptW^{(\ell)}_{j}:=\exp(\frac{1}{\ell}\varphi^{*}_{\omega}(\ell j)),\;j\in\mathbb{N}.$
@@
Since the definition of the weighted spaces and of $\mathcal{M}_{\omega}$ involves the conjugate $\varphi^{*}_{\omega}$ from (1.1) we comment now on properties of this conjugate for arbitrary $\omega:[0,+\infty)\rightarrow[0,+\infty)$ in detail; see also [7, Rem. 1.3, Lemma 1.5]. It turns out that even a more general setting is sufficient to introduce $\mathcal{M}_{\omega}$. Indeed, it suffices to assume that $\omega$ is non-decreasing and $\log(t)=o(\omega(t))$ as $t\rightarrow+\infty$ (i.e. $(\omega_{3})$) in order to ensure well-definedness of the Young conjugate and hence of each sequence $\mathbf{W}^{(\ell)}$; we refer also to the comments in [28, Sect. 3.1], [26, Sect. 2.7, (a)(c)(a)-(c)], [34, Sect. 6 $(III)$]. For future reference we call any non-decreasing function $\omega$ satisfying $(\omega_{3})$ matrix admissible.
@@
Summarizing, assume that $\omega$ is non-decreasing and such that $(\omega_{3})$ holds. Then $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$, so $\omega$ is a weight function in the notion of [12] (and [27], [26]), and by (a)(e)(a)-(e) it follows that each $\mathbf{W}^{(\ell)}$ is a well-defined sequence of non-negative real numbers, each $\mathbf{W}^{(\ell)}$ is log-convex (i.e. $(M.1)$ in [15] is valid), $\lim_{j\rightarrow+\infty}(W^{(\ell)}_{j})^{1/j}=+\infty$ for all $\ell>0$, and finally $W^{(\ell_{1})}_{j}\leq W^{(\ell_{2})}_{j}$ for all $j\in\mathbb{N}$, $0<\ell_{1}\leq\ell_{2}$. Thus basically “non-decreasing” and $(\omega_{3})$ for $\omega$ are sufficient to introduce $\mathcal{M}_{\omega}$ according to (2.13) and this explains the notion “matrix admissible”.
If $\omega$ is also normalized, then each $\mathbf{W}^{(\ell)}$ is normalized; i.e. $1=W^{(\ell)}_{0}\leq W^{(\ell)}_{1}$. Note that, when $\omega$ is not normalized, then one can always switch to an equivalent weight $\sigma$ which is normalized and this notion of equivalence transfers in a precise manner to the corresponding associated weight matrices and which ensures the equality of the corresponding (matrix) weighted spaces; we refer to [34, Sect. 6 $(III)$, $(6.6)$].
@@
Moreover, if $\omega$ is matrix admissible and satisfies in addition $(\omega_{1})$, then also the equalities $\mathcal{E}_{\{\mathcal{M}_{\omega}\}}=\mathcal{E}_{\{\omega\}}$ and $\mathcal{E}_{(\mathcal{M}_{\omega})}=\mathcal{E}_{(\omega)}$ (as l.c.v.s.) hold. Here the classes are introduced directly (by definition) via weighting the derivatives in terms of $\varphi^{*}_{\omega}$ (cf. [24, Sect. 4 & 5]). For this fact estimate [24, $(5.10)$] is required and for its proof $(\omega_{4})$ is not needed. The same comment applies to analogously defined weighted spaces.
If in addition $(\omega_{6})$ is valid, then by [24, Lemma 5.9, $(5.11)$] all $\mathbf{W}^{(\ell)}$ are equivalent; i.e. $0<\inf_{j\in\mathbb{N}_{>0}}(W^{(\ell_{1})}_{j}/W^{(\ell_{2})}_{j})^{1/j}\leq\sup_{j\in\mathbb{N}_{>0}}(W^{(\ell_{1})}_{j}/W^{(\ell_{2})}_{j})^{1/j}<+\infty$ for all $\ell_{1},\ell_{2}>0$. However, for the converse implication which is also shown in [24, Lemma 5.9] the double-conjugate and hence $(\omega_{4})$ is required. And also for the characterization in [24, Cor 5.8 (2)], which establishes a connection to condition “moderate growth” (i.e. $(M.2)$ from [15]) for some/each $\mathbf{W}^{(\ell)}$, the convexity is used by involving the equivalence between $\omega$ and (each) $\omega_{\mathbf{W}^{(\ell)}}$ (see [24, Lemma 5.7]).
@@
Recall for arbitrary functions $\sigma,\omega:[0,+\infty)\rightarrow[0,+\infty)$ the notation $\sigma\geq\omega$ which means $\sigma(t)\geq\omega(t)$ for all $t\geq 0$ (see Section 2.3).
@@ 3.1. Basic definitions
$\mathcal{L}^{\infty}_{\{\mathcal{W}\}}:=\bigcup_{\ell>0}\mathcal{L}^{\infty}_{\omega^{\ell}},\hskip 15.0pt\mathcal{L}^{\infty}_{(\mathcal{W})}:=\bigcap_{\ell>0}\mathcal{L}^{\infty}_{\omega^{\ell}},$
endowed with their natural topologies. Here, $\operatorname{ess}\sup$ denotes the essential supremum. Moreover, for any $1\leq p<\infty$ we assume that $\mathcal{W}$ is continuous and for all (parameters) $\ell>0$ let us set
@@
$\mathcal{L}^{p}_{\{\mathcal{W}\}}:=\bigcup_{\ell>0}\mathcal{L}^{p}_{\omega^{\ell}},\hskip 15.0pt\mathcal{L}^{p}_{(\mathcal{W})}:=\bigcap_{\ell>0}\mathcal{L}^{p}_{\omega^{\ell}}.$
@@ 3.2. On the unboundedness of weight function matrices
    The following are equivalent:
      Condition (3.2) holds,
      $L^{\infty}=\mathcal{L}^{\infty}_{\{\mathcal{W}\}}$ holds as sets,
      $L^{\infty}=\mathcal{L}^{\infty}_{\{\mathcal{W}\}}$ holds as l.c.v.s.
@@
    The following are equivalent:
      Condition (3.3) holds,
      $L^{\infty}=\mathcal{L}^{\infty}_{(\mathcal{W})}$ holds as sets,
      $L^{\infty}=\mathcal{L}^{\infty}_{(\mathcal{W})}$ holds as l.c.v.s.
  • $(iii)$
    If $\mathcal{W}$ is in addition continuous, then the following are equivalent:
      Condition (3.2) holds,
      $L^{p}=\mathcal{L}^{p}_{\{\mathcal{W}\}}$ holds as sets for some/any $1\leq p<\infty$,
      $L^{p}=\mathcal{L}^{p}_{\{\mathcal{W}\}}$ holds as l.c.v.s. for some/any $1\leq p<\infty$.
  • $(iv)$
    If $\mathcal{W}$ is in addition continuous, then the following are equivalent:
      Condition (3.3) holds,
      $L^{p}=\mathcal{L}^{p}_{(\mathcal{W})}$ holds as sets for some/any $1\leq p<\infty$,
      $L^{p}=\mathcal{L}^{p}_{(\mathcal{W})}$ holds as l.c.v.s. for some/any $1\leq p<\infty$.
@@ 3.3. On the non-triviality of weighted $L^{p}$-type spaces
Proof. $(i)$ If (3.3) is valid, then $\mathcal{L}^{\infty}_{(\mathcal{W})}=L^{\infty}\neq\{0\}$ by $(ii)$ in Theorem 3.3.
@@
$\psi^{\infty}_{\mathcal{W}}(x):=e^{-\sigma^{\infty}(x)^{2}}.$
So for any $\ell>0$ and all $x\in\mathbb{R}^{d}$ such that $|x|\in[n,n+1)$ one has $|\psi^{\infty}_{\mathcal{W}}(x)|e^{\omega^{\ell}(x)}=e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{2}}$. Let $\ell\geq\ell_{0}$ be fixed and hence for all $|x|\in[n,n+1)$ with $n+1\geq\ell$ we have $|\psi^{\infty}_{\mathcal{W}}(x)|e^{\omega^{\ell}(x)}\leq e^{\omega^{\ell}(x)(1-\omega^{\ell}(x))}$ and this last expression tends to 0 as $|x|\rightarrow+\infty$. Thus,
@@
$(ii)$ If (3.3) holds, then $\mathcal{L}^{p}_{(\mathcal{W})}=L^{p}\neq\{0\}$ by $(iv)$ in Theorem 3.3.
@@
Indeed, it suffices to take
$b_{n}:=\frac{\log(\omega^{n}(n+1))+\log(2)}{\log(\omega^{n}(n))}(>1).$
@@
$=\sum_{n<\ell}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{a_{n}b_{n}}}dx+\sum_{n\geq\ell}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{a_{n}b_{n}}}dx$
$\leq\sum_{n<\ell}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{a_{n}b_{n}}}dx+\sum_{n\geq\ell}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{n}(x)-(\omega^{n}(x))^{a_{n}b_{n}}}dx$
$\leq\sum_{n<\ell}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{a_{n}b_{n}}}dx+\sum_{n\geq\ell}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}\frac{1}{(1+|x|^{2})^{d+1}}dx<+\infty.$
@@ 4. Characterization of inclusion relations for weighted $L^{p}$-type spaces
Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ and $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be weight function matrices and assume that both matrices satisfy
(4.3) $\forall\;\ell>0\;\exists\;n>0\;\exists\;L\geq 1\;\forall\;x,y\in\mathbb{R}^{d}:\;\;\;\omega^{\ell}(x+y)\leq\omega^{n}(x)+\omega^{n}(y)+L;$
i.e. (3.10) for the same matrix. Then the following are equivalent:
@@
Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ and $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be weight function matrices and assume that both matrices satisfy
(4.4) $\forall\;n>0\;\exists\;\ell>0\;\exists\;L\geq 1\;\forall\;x,y\in\mathbb{R}^{d}:\;\;\;\omega^{\ell}(x+y)\leq\omega^{n}(x)+\omega^{n}(y)+L;$
i.e. (3.11) for the same matrix. Then the following are equivalent:
@@
Note: (4.3) resp. (4.4) yield the fact that all weighted spaces under consideration in the above results are translation invariant.
@@ 4.1. Proof of the main characterizing results
Proof. $(i)$ First, by (4.3) applied to the matrix $\mathcal{T}$ we get:
@@
Now, in view of Remark 4.4 we can assume that there exists $0\neq f\in\mathcal{L}^{p}_{(\mathcal{S})}$ and (4.3) for $\mathcal{S}$ gives via $(i)$ in Proposition 3.6 that $f_{x_{0}}\in\mathcal{L}^{p}_{(\mathcal{S})}$ for all $x_{0}\in\mathbb{R}^{d}$. The idea is to apply (4.5) to the family $\{f_{x_{0}}:x_{0}\in\mathbb{R}^{d}\}$.
Let $\ell>0$ be given, arbitrary but from now on fixed. Let $\ell^{\prime}>0$ be the index appearing in the estimate in the first step and being related to $\ell$ via (4.3) for the matrix $\mathcal{T}$. Then apply (4.5) to $\ell^{\prime}$ and let $\ell^{\prime\prime}$ be the index subject to $\ell^{\prime}$ in this property.
@@
The next result deals with the Roumieu case.
@@ 4.2. Alternative proofs for the Roumieu-type
(4.12) $\exists\;\ell^{\prime}_{0}>\ell_{0}>0:\;\;\;\liminf_{|x|\rightarrow+\infty}\omega^{\ell^{\prime}_{0}}(x)-\omega^{\ell_{0}}(x)>0,$
@@
    If (4.11), (4.12) holds for some indices $\ell^{\prime}_{0}>\ell_{0}$, then by the order of the elements in the matrix also for all $\ell^{\prime},\ell$ satisfying $\ell^{\prime}\geq\ell^{\prime}_{0}>\ell_{0}\geq\ell>0$.
@@
    One shall note that (4.11), (4.12) and (4.13) are crucial for the techniques in this section but they never can be satisfied for any simple weight function matrix $\mathcal{W}:=\{\omega\}$. On the other hand, (4.3) and (4.4) hold for any sub-additive $\omega$ (and coincide in the simple case) and hence the main results Theorems 4.1, 4.2 and Proposition 3.6 can be applied, too.
@@
$\|\theta^{p}_{\omega^{\ell}}\|^{p}_{p,\omega^{\ell^{\prime}}}=\int_{\mathbb{R}^{d}}e^{\omega^{\ell^{\prime}}(x)-\omega^{\ell}(x)}dx\geq\int_{\mathbb{R}^{d}}e^{\omega^{\ell^{\prime}_{0}}(x)-\omega^{\ell_{0}}(x)}dx=+\infty,$
and the last equality is valid since by assumption there exists some $\epsilon>0$ and $t_{\epsilon}>0$ such that $\omega^{\ell^{\prime}_{0}}(x)-\omega^{\ell_{0}}(x)\geq\epsilon$ for all $x\in\mathbb{R}^{d}$ satisfying $|x|\geq t_{\epsilon}$. However, in general it is not clear if already (4.11) implies this estimate; see Remark 4.12. ∎
@@
Note: This statement is more general than $(ii)$ in Proposition 4.6 since no stability under translation (i.e. condition (4.4)) is required.
@@
Proof. $(iii)\Rightarrow(i)$ follows immediately by definition of the spaces and $(i)\Rightarrow(ii)$ is clear. For $(ii)\Rightarrow(iii)$ follow the proof of Proposition 4.9 and replace $\exists\;\ell^{\prime}>0$” by $\forall\;\ell^{\prime}>0$”. ∎
@@
    (4.4) obviously implies (4.16) and (4.3) implies (4.17).
@@
thus $\mathcal{S}\{\preceq\}\mathcal{T}$ is verified between the indices $\ell^{\prime\prime}(\geq\ell)$ and $\ell_{1}$.
@@
And for $x\in\mathbb{R}^{d}$ with $0\leq|x|<1$ one has $\tau^{\ell^{\prime\prime}_{1}}(x)\leq\tau^{\ell^{\prime\prime}_{1}}(1)\leq\sigma^{\ell}(x)+A$ with $A:=\tau^{\ell^{\prime\prime}_{1}}(1)$. Summarizing, again $\mathcal{S}\{\preceq\}\mathcal{T}$ is verified between the indices $\ell$ and $\ell^{\prime\prime}_{1}(\leq\ell_{1})$. ∎
@@
Without the additional assumptions on the weight matrices it is not clear that (4.18) implies $\lim_{|x|\rightarrow+\infty}e^{\tau^{\ell_{1}}(x)-\sigma^{\ell}(x)}=0$ or, what is enough to conclude, $\sup_{x\in\mathbb{R}^{d}}e^{\tau^{\ell_{1}}(x)-\sigma^{\ell}(x)}<+\infty$. This is due to the fact that in general $t\mapsto e^{\tau^{\ell_{1}}(t)-\sigma^{\ell}(t)}$ might behave very irregular: This function can have (infinitely many) “exploding peaks” but being sufficiently small “nearly everywhere on $\mathbb{R}$” and hence (4.18) still can be valid. In view of Remark 4.7 this comment becomes relevant in the simple case when no further index is appearing.